Aqueous sulfuric acid H2SO4 will react with solid sodium hydroxide NaOH to produce aqueous sodium sulfate Na2SO4 and liquid water H2O. Suppose 5.88 g of sulfuric acid is mixed with 6.2 g of sodium hydroxide. Calculate the maximum mass of sodium sulfate that could be produced by the chemical reaction. Be sure your answer has the correct number of significant digits.
1
Expert's answer
2019-09-30T06:19:51-0400

Reaction between them is given by-

"2NaOH\ \ +H_2SO_4\ \to\ Na_2SO_4\ \ +2H_2O"


Molar mass of "H_2SO_4=98\ g\ mole^{-1}"

Moles of sulphuric acid taken = "\frac{5.88}{98}=0.06\ mole"

Moles of sodium hydroxide taken = "\frac{6.2}{40}=0.155\ mole"

here sulphuric acid is limiting agent so

reaction will proceed according to sulphuric acid

"0.06\ mole\ H_SO_4\ will \ produce\ same\ amount\ of \ \ Na_2SO_4\\"

Maas of sodium sulphate produced = "moles\times molar\ mass=0.06\times 142=8.52\ g"